# Interactive Learning of Physical Object Properties Through Robot Manipulation and Database of Object Measurements

### v1

@@ V Database of object properties and measurements for robot manipulation
Object datasets related to robot manipulation typically contain object models and visual-based representations like images, point clouds, or meshes (e.g., YCB [12]). Datasets containing physical object properties are an exception (e.g., [17, 18, 19]).
The measurements collected in this work provided the foundation for a database of measurements of object physical properties. These measurements are naturally dependent on the objects manipulated and also highly dependent on the embodiment of the robot (e.g., stiffness measurements on the gripper used and the force/effort feedback)—unlike for images where the camera that recorded the images plays less critical role. To be useful, such a database needs to cover different setups: different robots, grippers, sensors, etc. and their description in the metadata. The structure of the database we created is shown in Fig. 9; it is available at https://cmp.felk.cvut.cz/ipalm/.
@@ VI Conclusion, Discussion, Future Work
We presented an algorithmic framework and experimental setup for learning physical object properties through robot manipulation. Our approach involved exploratory action selection to maximize IG about objects on a table in front of a robot manipulator. At its core, the framework employs a BN model to represent the probabilistic relationships between different object properties. The BN consists of nodes corresponding to various object properties, including categorical properties like Category and Material, as well as continuous properties such as Elasticity, Density, and Volume. We demonstrated the effectiveness of our method through experimental evaluation, showcasing its ability to automatically extract properties such as material composition, mass, volume, and stiffness. The algorithm outperformed random action selection baseline and showed promising results in discovering object properties on a partially adversarial object dataset (with “fake objects” like a stiff sponge). The algorithm can automatically terminate based on the expected IG of the available actions. The code used in this work is available at https://github.com/ctu-vras/actsel. The object database and source code are available at https://cmp.felk.cvut.cz/ipalm/.
@@
Another source of information degradation is the inference of categorical weights from the continuous GMMs using the EM algorithm. This stems from uncertainties on the numerical level when the belief distribution mean moved too far from the reference labels. Therefore, the only information that the EM algorithm possesses at the location of reference labels is from the far tails of the normal distributions. In the process of normalization to unit-sum probabilities from very small numbers, the numerical errors provide a large variance in results. We tried to tackle this problem using Monte Carlo Markov Chains, more specifically Gibbs’ sampling method, although this has shown to be too computationally expensive to run in real-time.
As part of future work, community engagement and dataset expansion are essential endeavors. By making our dataset and codebase openly accessible, we aim to foster collaboration and knowledge-sharing among researchers in the field of robot learning and manipulation. We encourage contributions from the community to enrich the dataset with diverse object measurements and annotations, facilitating a broader range of research applications and algorithm evaluations. Additionally, active engagement with the community allows us to gather valuable feedback and insights, guiding the evolution of our framework and ensuring its relevance and usefulness in real-world settings. Together, these efforts contribute to the collective advancement of robot learning and manipulation capabilities, ultimately driving innovation and progress in the field.
Another promising avenue for advancement lies in the development of adaptive exploration strategies. By leveraging techniques from reinforcement learning and advanced decision-making algorithms, robots can autonomously learn and refine their exploration policies over time. Adaptive exploration strategies can dynamically adjust the selection of actions based on the robot’s past experience and the current state of uncertainty in object properties. This allows robots to prioritize actions that are more likely to yield informative measurements, leading to more efficient and effective learning of physical object properties.

### v2

@@ V Database of object properties and measurements for robot manipulation
Object datasets related to robot manipulation typically contain object models and visual representations like images, point clouds, or meshes (e.g., YCB [15]). Datasets containing physical object properties are an exception (e.g., [20, 21, 22]).
The measurements collected in this work provided the foundation for a database of measurements of object physical properties. These measurements are naturally dependent on the objects manipulated and also highly dependent on the embodiment of the robot (e.g., stiffness measurements on the gripper used and the force/effort feedback)—unlike for images where the camera that recorded the images plays a less critical role. To be useful, such a database needs to cover different setups: different robots, grippers, sensors, etc. and their description in the metadata. The structure of the database we created is shown in Fig. 9; it is available at https://cmp.felk.cvut.cz/ipalm/.
@@ VI Conclusion, Discussion, Future Work
We presented an algorithmic framework and experimental setup for learning physical object properties through robot manipulation. Our approach involved exploratory action selection to maximize expected Information Gain (IG) about objects on a table in front of a robot manipulator. At its core, the framework employs a Bayesian Network (BN) model to represent the probabilistic relationships between different object properties. The BN consists of nodes corresponding to various object properties, including categorical properties like Category and Material, as well as continuous properties such as Elasticity, Density, and Volume. We demonstrated the effectiveness of our method through experimental evaluation, showcasing its ability to automatically extract properties such as material composition, mass, volume, and stiffness. The algorithm outperformed random action selection baseline and showed promising results in discovering object properties on a partially adversarial object dataset (with “fake objects” like a stiff sponge). The algorithm can automatically terminate based on the expected IG of the available actions. The code used in this work is available at https://github.com/ctu-vras/actsel. The object database and source code are available at https://cmp.felk.cvut.cz/ipalm/.
@@
Another source of information degradation is the inference of categorical weights from the continuous GMMs using the EM algorithm. This stems from uncertainties on the numerical level when the belief distribution mean moved too far from the reference labels. Therefore, the only information that the EM algorithm possesses at the location of reference labels is from the far tails of the normal distributions. In the process of normalization to unit-sum probabilities from very small numbers, the numerical errors provide a large variance in results. We tried to tackle this problem using Monte Carlo Markov Chains, more specifically Gibbs’ sampling method, although this has shown to be too computationally expensive to run in real time.
As part of future work, community engagement and dataset expansion are essential endeavors. By making our dataset and codebase openly accessible, we aim to foster collaboration and knowledge sharing among researchers. We encourage contributions from the community to enrich the dataset with diverse object measurements and annotations, facilitating a broader range of research applications and algorithm evaluations. Additionally, active engagement with the community allows us to gather valuable feedback and insights, guiding the evolution of our framework and ensuring its relevance and usefulness in real-world settings. Together, these efforts contribute to the collective advancement of robot learning and manipulation capabilities, ultimately driving innovation and progress in the field.
Another promising avenue for advancement lies in the development of adaptive exploration strategies. By leveraging techniques from reinforcement learning and advanced decision-making algorithms, robots can autonomously learn and refine their exploration policies over time. Adaptive exploration strategies can dynamically adjust the selection of actions based on the past experience and the current state of uncertainty in object properties. This allows robots to prioritize actions that are more likely to yield informative measurements, leading to more efficient and effective learning of physical object properties. An expert policy will be a more challenging baseline than random action selection.